OpenStudy (anonymous):

slope of a line passing through the point (4,-2) and perpendicular to the line y=4x +2?

OpenStudy (amistre64):

lines tend to be defined by a point and a slope: usually of the form: y = mx -mPx + Py where m is the given slope and the point (Px,Py) is on the line

OpenStudy (amistre64):

perp lines have the property that their slopes multiply to get -1 m1 * m2 = -1 ; which means they are negative reciprocals of each other

OpenStudy (amistre64):

your m1 = 4 here, so your perp line needs m=-1/4 and your already given a point for the line to plug in

OpenStudy (amistre64):

try to form the eq from that and ill let you know how good it is if you need more help with it